Calculate the number of unique 4 person groups that can be formed by selecting from eight eligible candidates.
Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(8,4) = (8!)/(4!*(8-4)!)

C(8,4) = (8!)/(4!*4!)

C(8,4) = (8*7*6*5*4!)/(4!*4!)

C(8,4) = (8*7*6*5)/(4!)

C(8,4) = (8*7*6*5)/(4*3*2*1)

C(8,4) = (1680)/(24)

C(8,4) = 70

So there are 70 ways to do this assuming order doesn't matter.